Klapperich Lab:Notebook/Lab Meeting Notes/2008/08/25

From OpenWetWare
Jump to: navigation, search

25 August 2008 Lab Meeting Agenda

† Announcements

  • Discussion of Time for Fall Semester Group Meetings

Type when you are NOT available for group meetings HERE:

Cathie: MW afternoon, F all day
Sonali, Jaephil, MinCheol, open all week
Qingqing T/TH 10-12, 2-4pm, M 12-3, one lab still deciding
Frank M 8-10 & 12-1, T/R 11-6:30, W/F 9-3
Jane TF 3-5 (designated SEM time)
Hussam MW 10-12pm , T TH 4-6pm F:10-1pm
Mark MW 10-12 & 2-4; T Th 4-6; F 10-1

† Other Announcements

  • Cathie Bring IBM laptop and drive from home
  • No meeting 9/1.
  • Register for CIMIT conference if you want. Irene can reimburse.
  • Register for IMECE in Boston
  • Register for MRS Boston if you want to.

† Flu R01

  • Meeting set F afternoon (8/29)
  • Planning plaque assay - protocol decided

† SEPSIS Project/Chip Bonding Issues

  • Check if Anlee can do backscatter electrons on new SEM
  • EF experiments results later today
  • Chips did not bond - that were made last weekend.
  • Controlled cooling with air plasma bond okay with Zeonor plaque.
  • Topas COC resin may be better to try. Try and buy some. (this is what the plaques are made of) The COC might shrink less. (co-polymer).
  • Jaephil will help Sonali get some new material for the sepsis chips.
  • Wash is acetone then methanol sonication + DI water sonication. SEE UPDATED PROTOCOL!

† COBRA

  • Jaephil and Mark making humidity and flowrate control box (temp too) for the concentration system.
  • IEEE abstract submitted on 8/24.
  • IBC for our lab STILL pending - Alexis will give us the requested revisions
  • Cathie will submit this week (8/29)
  • Tests on chips with covers underway
  • Jane is investigating the multilayer design

† Valve Array

  • Cathie contact Sumita Pennarthur at UCSB about her valves
  • Waiting for fixture from Melissa
  • Mark's Part is complete at this point
  • Basic software in place

† Fraunhofer: LOAC

  • Still troubleshooting

† Biointerfaces group

  • Brett and Mincheol tried the SU8 process.
  • Hussam made 6.5 microns SU8 - Using the Hard contact method
  • Hussam made the PDMS part. Correct part, but overexposed. The thickness is too thin
  • Mincheol will try the device as we have it
  • Hussam will make 20 micron high with 1 micron space this week/or Brett
  • SU8 thickness is important step in this process. So, ALIQUOT SU8 when it arrives.
  • Solvent evaporation from the SU8 can be a real problem.
  • Use the same spinner every time for this project.
  • Wafer cleaning should be the same every time, too
  • Hussam, Brett and MinCheol are going to make a protocol for this process
  • AZP 4620 from Clariant. Hussam will try this photoresist (positive)

† CIMIT

  • Meet with Sepsis Nurse Tuesday.
  • QQ submitted abstract.
  • Jane submited.

† PCR

  • QQ Model has 30 cycles. MinCheol will try the model. Need to upgrade the Star-CD program. Dan Kamalic has to do this.
  • Emailed Dan today.
  • QQ passed qual!! yeah!!!
  • Old heaters worked 20+ times, but then are not homogenous across chip.
  • Waiting for new heaters to try again. Other brand was not good. Will use copper in the meantime for temp control.
  • PUMP problem. Speed control at low flow rates is not stable.
  • What are the specs on Wong's pump?
  • What are the specs on our pumps?
  • Can KDS calibrate or fix it?
  • We need at least one more pump that does multiple channels. Mincheol will check the specs?
  • What requirements does everyone have for their "IDEAL" pump.

† RCA

  • STILL Working on a test chip to examine mat'ls compatibility
  • 50 microliter test chip made/ may be and issue getting the stuff out. Especially after the wash.
  • RCA meeting with Heiko, Jaephil, Sonali this week.
  • Whitepaper composition.

† Senior project

  • Megan still gone.
  • New students will be recruited.

† F31: Cochlea

  • Nathan in Oregon

† Silica Optimization (Lambda):

  • Ongoing

Hussam's Update